OpenStudy (anonymous):

Can somebody please show me the steps needed: f(x) = 1 / sqrt(2x - 1) 1) Use the definition of a derivative to find the gradient of the tangent to f(x) at any point x 2) Hence determine the value of the gradient at x = 1

OpenStudy (cj49):

f(x) = √(1+2x) f(x+h) = √(1+2x+2h) f(x+h) - f(x) = √(1+2x+2h) - √(1+2x) multiply and divide RHS by √(1+2x+2h) + √(1+2x) f(x+h) - f(x) = [ (1+2x+2h) - (1+2x) ] /[√(1+2x+2h) +√(1+2x) ] = (2h) /[√(1+2x+2h) +√(1+2x) ] [f(x+h) - f(x) ] /h = 2 /[√(1+2x+2h) +√(1+2x) ] when h--> 0 The expression reduces to 2 /[√(1+2x) +√(1+2x) ] f ' (x) = 2/2√(1+2x) = 1/√(1+2x)

OpenStudy (precal):

now just do f '(1) and you have all of your answers

OpenStudy (anonymous):

Where did you get the f(x) = √(1+2x) from?

OpenStudy (precal):

that is f ' (x) not f (x)

OpenStudy (anonymous):

Thank you! That clears up a bit of confusion... Could you maybe show me the steps to determine the derivative of \[f(x) = \frac{ 1 }{ \sqrt{2x -1} }\]

OpenStudy (precal):

cj49 already did

OpenStudy (precal):

cj49 did it by definition of a derivative

OpenStudy (precal):

you could verify it by using chain rule, but that was not your instructions

OpenStudy (anonymous):

Sorry, I'm still really confused... Can you correct me where I'm wrong? \[f'(x) = \lim_{h \rightarrow 0} \frac{ f(x + h) - f(x) }{ h }\] \[f'(x) = \lim_{h \rightarrow 0} \frac{ \frac{ 1 }{ \sqrt{2(x + h) - 1} } - \frac{ 1 }{ \sqrt{2x - 1} } }{ h }\] And from there?

OpenStudy (anonymous):

I'm just not entirely sure on how to find the derivative...

OpenStudy (precal):

you have to multiply the top and bottom by the conjugate

OpenStudy (precal):

multiply and divide RHS by √(1+2x+2h) + √(1+2x) this step right here is where cj49 did so

OpenStudy (anonymous):

So... \[f'\left( x \right) = \lim_{h \rightarrow 0} \frac{ \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } - \frac{ 1 }{ \sqrt{2x - 1} }}{ h } \times \frac{ \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} }}{ \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} } }\] then \[f'\left( x \right) = \lim_{h \rightarrow 0} \frac{ \frac{ 1 }{ 2\left( x + h \right) - 1 } - \frac{ 1 }{ 2x - 1}}{ \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} } }\]?

OpenStudy (precal):

what happened to the h on the bottom?

OpenStudy (anonymous):

I forgot about that... \[f'\left( x \right) = \lim_{h \rightarrow 0} \frac{ \frac{ 1 }{ 2\left( x + h \right) - 1 } - \frac{ 1 }{ 2x - 1}}{ h\left( \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} } \right)}\] So this?

OpenStudy (precal):

ok now distribute that 2(x+h)

OpenStudy (precal):

on the top

OpenStudy (precal):

then get common denominators for the top part

OpenStudy (anonymous):

So \[\frac{ 1 }{ 2\left( x + h \right) - 1} - \frac{ 1 }{ 2x -1 }\] becomes \[\frac{ 1 }{ 2x + 2h - 1} - \frac{ 1 }{ 2x -1 }\] and then \[\frac{ \left( 2x - 1 \right) - \left( 2x + 2h - 1 \right) }{ \left( 2x + 2h - 1 \right)\left( 2x - 1 \right)}\]?

OpenStudy (anonymous):

And then\[\frac{ -2h }{ 4x ^{2} + 4xh -4x -2h + 1 }\]?

OpenStudy (precal):

sorry I ran off for a bit, let me look at this

OpenStudy (precal):

ok that is just the top correct?

OpenStudy (anonymous):

Just the top part

OpenStudy (precal):

ok let me get out some paper, somehow I think I can be more simpler. we might have made it look complex

OpenStudy (precal):

ok give me a moment to type it out so recall that when you divide fractions, you leave the top alone, change division to multiplication and flip the second fraction, correct?

OpenStudy (precal):

\[\frac{ -2h }{ (2x+2h-1)(2x-1) }\left[ \left( h \right)\frac{ 1 }{ \sqrt{2(x+h)-1} } +\frac{ 1 }{ \sqrt{2x-1} }\right]\]

OpenStudy (precal):

ok boy did I mess that up, that h needs to be a denominator

OpenStudy (precal):

basically it is going to cancel out with the h from -2h do you see it?

OpenStudy (precal):

your next step is to sub h=0 that represents as h approaches 0

OpenStudy (precal):

do you understand or should I try to type it again.....

OpenStudy (precal):

\[\frac{ -2h }{ 2(x+h)-1 }\left[ \frac{ 1 }{ h \sqrt{2(x+h)-1} }+\frac{ 1 }{ \sqrt{2x-1} } \right]\]

OpenStudy (precal):

ok like this, now do you see that the h from -2h and h cancel out

OpenStudy (precal):

\[\frac{ -2 }{ 2(x+h)-1 }\left[ \frac{ 1 }{ \sqrt{2(x+h)-1} }+\frac{ 1 }{ \sqrt{2x-1} } \right]\]

OpenStudy (precal):

\[\frac{ -2 }{ 2(x+0)-1 }\left[ \frac{ 1 }{ \sqrt{2(x+0)-1} }+\frac{ 1 }{ \sqrt{2x-1} } \right]\]

OpenStudy (precal):

\[\frac{ -2 }{ 2x-1 }\left[ \frac{ 1 }{ \sqrt{2x-1} }+\frac{ 1 }{ \sqrt{2x-1} } \right]\]

OpenStudy (precal):

\[\frac{ -2 }{ 2x-1 }\left[ \frac{ 2 }{ \sqrt{2x-1} } \right]\]

OpenStudy (precal):

I almost have it, I know the answer ok so do you agree 2x-1 is to the first power (2x-1)^1 so when you multiply across the denominator is (2x-1)^(3/2)

OpenStudy (precal):

numerator is -1

OpenStudy (precal):

so I know I mess up somehow......

OpenStudy (precal):

\[\frac{ -1 }{ (2x-1)^{3/2} }\]

OpenStudy (precal):

this is your answer..........

OpenStudy (anonymous):

So then\[\frac{ -4 }{ \left( 2x - 1 \right) \sqrt{2x -1}}\]becomes\[\frac{ -4 }{ \left( 2x - 1 \right)^{\frac{ 3 }{ 2 }}}\]? I kinda got lost near the end there... How did you get the -1 for the numerator?

OpenStudy (precal):

that is the point, I did mess up since it should be -1, I know this because I used the chain rule......

OpenStudy (precal):

let me redo the entire problem..........having more problems typing it in rather than doing it

OpenStudy (precal):

\[\frac{ -2h }{ 2(x+h)-1 }\left[ \frac{ 1 }{ h \sqrt{2(x+h)-1} }+\frac{ 1 }{ \sqrt{2x-1} } \right]\]

OpenStudy (anonymous):

Cancel out the 2's over there?

OpenStudy (precal):

I mess up on the second part....... because I need to simplify everything in brackets no you can not cancel out the 2's

OpenStudy (precal):

\[h\left[ \frac{ 1 }{ \sqrt{2(x+h)-1} }+\frac{ 1 }{ \sqrt{2x-1} } \right]\]

OpenStudy (anonymous):

I forgot about the -1 in the denominator... whoops

OpenStudy (precal):

lets simplify this first......

OpenStudy (precal):

\[\frac{ h \left( \sqrt{2x-1}+\sqrt{2(x+h)-1} \right) }{ (\sqrt{2(x+h)-1})(\sqrt{2x-1}) }\]

OpenStudy (precal):

\[\frac{ -2h }{ 2(x+h)-1 }/\frac{ h \left( \sqrt{2x-1}+\sqrt{2(x+h)-1} \right) }{ (\sqrt{2(x+h)-1})(\sqrt{2x-1}) }\]

OpenStudy (precal):

remember this is divided by now we need to change it to times and flip it over, ok wish me luck

OpenStudy (anonymous):

Good luck! :P

OpenStudy (precal):

\[\frac{ -2h }{ 2(x+h)-1 }\left[ \frac{ (\sqrt{2(x+h)-1})(\sqrt{2x-1}) }{ h( \sqrt{2x-1}+\sqrt{2(x+h)-1}} \right]\]

OpenStudy (precal):

\[\frac{ -2 }{ 2(x+h)-1 }\left[ \frac{ (\sqrt{2(x+h)-1})(\sqrt{2x-1}) }{ ( \sqrt{2x-1}+\sqrt{2(x+h)-1}} \right]\]

OpenStudy (precal):

this is the h cancels out now lets sub h=0 this is where h approaches 0

OpenStudy (precal):

\[\frac{ -2 }{ 2(x+0)-1 }\left[ \frac{ (\sqrt{2(x+0)-1})(\sqrt{2x-1}) }{ ( \sqrt{2x-1}+\sqrt{2(x+0)-1}} \right]\]

OpenStudy (precal):

\[\frac{ -2 }{ 2x-1 }\left[ \frac{ (\sqrt{2x-1})(\sqrt{2x-1}) }{ ( \sqrt{2x-1}+\sqrt{2x-1}} \right]\]

OpenStudy (precal):

\[\frac{ -2 }{ 2x-1 }\left[ \frac{ 2x-1 }{ ( \sqrt{2x-1}+\sqrt{2x-1}} \right]\]

OpenStudy (anonymous):

\[\frac{ -2 }{ 2x - 1 }\left[ \frac{ 2x - 1 }{ \sqrt{2x - 1} + \sqrt{2x - 1}} \right]\] \[\frac{ -2 }{ 2x - 1 }\left[ \frac{ 2x - 1 }{ 2\sqrt{2x - 1}} \right]\] \[\frac{ -1 }{ 1 }\left[ \frac{ 1 }{ \sqrt{2x - 1}} \right]\] \[\frac{ -1 }{\sqrt{2x - 1}}\]

OpenStudy (precal):

no but that is wrong because the bottom has to be to the 3/2 power

OpenStudy (precal):

@ganeshie8 @SolomonZelman Can you help me figure out how I messed up? I feel bad because I am trying to help someone and I have mess this up twice...

OpenStudy (anonymous):

I'm off to have supper quickly but I'll be right back... Thank you for trying to help me and good luck solving it!

OpenStudy (precal):

Thanks @ganeshie8 for taking a look at this problem

ganeshie8 (ganeshie8):

\[\large \begin{align} \\ f'\left( x \right) &= \lim_{h \rightarrow 0} \frac{ \frac{ 1 }{ 2\left( x + h \right) - 1 } - \frac{ 1 }{ 2x - 1}}{ h\left( \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} } \right)} \end{align} \]

ganeshie8 (ganeshie8):

in the next step, the denominator was multiplied w/o flipping... thats the only mistake i see... other than that, everything looks good to me !

ganeshie8 (ganeshie8):

\[\large \begin{align} \\ f'\left( x \right) &= \lim_{h \rightarrow 0} \frac{ \frac{ 1 }{ 2\left( x + h \right) - 1 } - \frac{ 1 }{ 2x - 1}}{ h\left( \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} } \right)} \\~\\ &= \lim_{h \rightarrow 0} \frac{ \frac{ -2h }{ (2\left( x + h \right) - 1)(2x-1) } }{ h\left( \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} } \right)} \\ \end{align} \]

ganeshie8 (ganeshie8):

\[\large \begin{align} &= \lim_{h \rightarrow 0} \frac{-2h }{ h(2\left( x + h \right) - 1)(2x-1)\left( \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} } \right)} \\ \end{align} \]

ganeshie8 (ganeshie8):

\[\large \begin{align} &= \lim_{h \rightarrow 0} \frac{-2 }{ (2\left( x + h \right) - 1)(2x-1)\left( \frac{ 1 }{ \sqrt{2\left( x + h \right) - 1} } + \frac{ 1 }{ \sqrt{2x - 1} } \right)} \\ \end{align} \]

ganeshie8 (ganeshie8):

guess we can take the limit as h is gone

OpenStudy (precal):

thanks so much........

ganeshie8 (ganeshie8):

np :)

OpenStudy (precal):

\[\frac{ -2 }{ (2x-1)(2x-1)\left[ \frac{ 1 }{ \sqrt{2x-1 }}+\frac{ 1 }{ \sqrt{2x-1} } \right] }\]

OpenStudy (precal):

\[\frac{ -2 }{ (2x-1)^{2}\left[ \frac{ 2 }{ \sqrt{2x-1 }} \right] }\]

OpenStudy (precal):

\[\frac{ -2 }{ \left[ \frac{ 2 (2x-1)^{2}}{ \sqrt{2x-1 }} \right] }\]

OpenStudy (precal):

\[\frac{ -2 }{ 2(2x-1)^{3/2} }\]

OpenStudy (precal):

\[\frac{ -1 }{ (2x-1)^{3/2} }\]

OpenStudy (precal):

finally the correct solution

OpenStudy (anonymous):

Thank you so much for helping me! I owe you a lot...

OpenStudy (precal):

no problem ganeshie8 really pointed out my mistake......